# From Intuition to Expertise: Rubric-Based Cognitive Calibration for Human Detection of LLM-Generated Korean Text

### v1

@@ Abstract
Distinguishing human-written Korean text from fluent LLM outputs remains difficult even for linguistically trained readers, who can over-trust surface well-formedness. We study whether expert detection can be treated as a learnable skill and improved through structured calibration. We introduce LREAD, a rubric derived from national Korean writing standards and adapted to target micro-level artifacts (e.g., punctuation optionality, spacing behavior, and register shifts). In a three-phase longitudinal blind protocol with Korean linguistics majors, Phase 1 measures intuition-only detection, Phase 2 enforces criterion-level scoring with explicit justifications, and Phase 3 evaluates domain-focused mastery on held-out elementary essays. Across phases, majority-vote accuracy increases from 60% to 100%, accompanied by stronger inter-annotator agreement (Fleiss’ $\kappa$: -0.09 $\rightarrow$ 0.82). Compared to state-of-the-art LLM detectors, calibrated humans rely more on language-specific micro-diagnostics that are not well captured by coarse discourse priors. Our findings suggest that rubric-scaffolded expert judgment can serve as an interpretable complement to automated detectors for non-English settings, and we release the full rubric and a taxonomy of calibrated detection signatures.
From Intuition to Expertise: Rubric-Based Cognitive Calibration for Human Detection of LLM-Generated Korean Text
@@ 1 Introduction
The proliferation of large language models (LLMs) has reshaped digital communication and intensified concerns about academic (perkins2023academic; peng-etal-2023-hidding; zhu2024embracing) and creative integrity (yang-etal-2024-survey; wu-etal-2025-survey). Recent models also generate text that is highly fluent and norm-conforming, which can appear authoritative even to trained readers. Although prior work reports systematic differences between human and machine writing, such as increased dependency depth in machine-generated outputs alongside greater semantic variety in human text (zanotto-aroyehun-2025-linguistic), these cues do not reliably translate into stable human judgments. In practice, human annotators often over-rely on surface well-formedness, as polished style and local grammaticality can obscure subtle structural irregularities, resulting in what we characterize as a persistent fluency trap for linguistically trained yet task-uncalibrated readers.
@@
A common expectation is that domain expertise should reduce this vulnerability. Evidence from English settings supports this view: frequent LLM users can outperform commercial systems such as GPTZero [^1] https://gptzero.me/ when identifying synthetic articles (russell-etal-2025-people). At the same time, the majority of detection evidence and tooling remains English-centered, while high-stakes misuse increasingly occurs across diverse languages and writing norms. Non-English settings introduce additional sources of ambiguity, including language-specific morphology, orthography, and punctuation conventions, which can alter both the appearance of fluency and the distribution of artifacts. These factors motivate language-specific, human-centered detection studies that characterize expert reasoning in practice and promote consistent and explainable judgments.
We focus on Korean as a representative case of a morphologically rich language with flexible spacing conventions. Within the Korean linguistic context, prior work identifies specific morphological and punctuational artifacts, including distinctive comma usage patterns, that distinguish LLM-generated Korean text from human-written Korean text (park-etal-2025-katfishnet). Our preliminary findings reveal a distinct failure mode for intuitive detection in this setting. Even Korean linguistics majors, despite formal training in grammar, discourse, and style, achieve only 60% accuracy in a blind attribution task. Their decisions further exhibit low initial inter-annotator agreement, suggesting not only the presence of expert error but also the absence of a shared evaluative criterion. This pattern aligns with the broader observation that humans struggle to draw stable boundaries between human- and machine-authored text, yet can improve through iterative feedback and structured intervention (dugan2023real). We therefore hypothesize that the primary bottleneck is the absence of an explicit and transferable inferential procedure, rather than an inherent deficit in linguistic competence.
We introduce the linguistically-regularized expert AI detection rubric (LREAD), a diagnostic framework derived from national linguistic standards for systematic textual analysis. LREAD functions as a cognitive scaffold that externalizes expert intuition into an explicit, criterion-level decision protocol requiring concrete evidence for each judgment. We integrate LREAD into a three-phase longitudinal blind test that traces the evolution of expert detection from intuition-driven screening to disciplined micro-diagnostics. Across phases, annotators improve detection accuracy from 60% to 100%. This improvement coincides with convergence in judgment, reflected in a marked increase in inter-annotator agreement measured by Fleiss’ $\kappa$. Calibrated experts also converge on a shared lexicon of LLM-generation cues, including cross-lingual interference (e.g., translationese) and model-specific stylistic accents.
@@ 2.1 Experimental Setup
We utilize the KatFish corpus (park-etal-2025-katfishnet), the first benchmark dataset specifically designed for detecting LLM-generated Korean text. We select argumentative essays as the primary genre for evaluation because they require a complex balance of logical reasoning and stylistic consistency. Our evaluation set comprises 30 samples selected through stratified random sampling from elementary, middle, and high school levels. Each level contains ten essays, consisting of two human-authored pieces and eight LLM-generated outputs. The LLM-generated portion includes two samples each from GPT-4o, Solar-10.7B, Qwen2-72B, and Llama3.1-70B.
We recruit three undergraduate students majoring in Korean Language and Literature to serve as human annotators. These annotators demonstrate high linguistic sensitivity while having minimal prior exposure to diverse large language models. During the initial blind test phase, we intentionally refrain from providing any formalized rubric or explicit decision criteria, allowing annotators to rely solely on their intuitive linguistic judgment. Each annotator is presented with one essay at a time and is instructed to determine whether the text was authored by a human or generated by an AI system.
For the automated baseline, we employ three state-of-the-art LLMs: GPT-5.2 Thinking, Gemini 3 Flash, and Claude Sonnet 4.5. We provide all participants with the same essay samples and instructions to perform a blind classification task without prior information regarding label distribution or source models.
@@
2.2 Human-LLM Detection Gap
We present the comparative results of the initial blind test in Table 1. Our findings reveal a substantial gap between intuition-driven human attribution and LLM-based reviewing under identical instructions. On this benchmark, LLM detectors achieve comparatively high zero-shot accuracy, whereas human linguistics experts are less reliable under intuition-only attribution.
We identify a 36.67%p gap between the human and LLM majority votes. These results demonstrate that the superficial fluency of modern LLMs successfully bypasses the intuitive linguistic filters of human experts. We find that linguistic expertise does not inherently confer robustness to surface-plausibility bias; experts can conflate technical well-formedness with human authorship and underestimate model-specific artifacts.
@@ 2.3 Analysis: Initial Human Perception
We summarize the heuristics that human experts use in the initial blind test to distinguish LLM-generated essays from human writing. These observations reflect the subjective filters that annotators apply before they access ground-truth labels or the formal rubric. We organize these heuristics around structural regularity and discourse plausibility.
@@
Annotators treat structural rigidity and mechanical consistency as primary signals of machine generation. They also maintain unusually regular surface patterns, including repetitive sentence openings and overly uniform punctuation, and they exhibit formatting artifacts in mixed-script contexts. Annotators further report that model-written essays adhere to the prompt with excessive literalness, yielding arguments that are normatively safe yet substantively shallow, and thereby constraining perspective-taking. Annotators also interpret conservative lexical choices and low semantic risk-taking as consistent with machine-generated text.
In contrast, annotators associate human authorship with linguistic fallibility and creative deviation. Human essays exhibit minor orthographic errors, localized grammatical inconsistencies, and idiosyncratic metaphors, yielding prose that is less polished yet more individual in voice. Annotators also perceive a more organic rhetorical progression in human writing, including personal anecdotes and emotionally grounded appeals. Overall, human annotators initially equate technical polish with artificiality and treat controlled inconsistency as a cue of human identity.
@@ 3 LREAD: A Human-Centered Cognitive Calibration Framework
3.1 Rubric Design from Linguistic Standards
We derive the LREAD rubric from the 2025 standards for argumentative writing proficiency issued by the National Institute of Korean Language. While these national standards provide a robust foundation for general writing assessment, we reconfigure the criteria to heighten sensitivity toward computational artifacts. Importantly, we design the rubric after the completion of the Phase 1 blind evaluation, once the three expert annotators were informed of the ground-truth labels. This post-hoc design choice ensures that the rubric systematically encodes the diagnostic cues that human experts identified during unconstrained judgment. We transform the standard assessment into a specialized detection tool by implementing two fundamental modifications.
@@ 3.1 Rubric Design from Linguistic Standards
Phase 1 Intuitive
Zero-shot detection task using
unstructured linguistic intuition.
Quantifying baseline susceptibility
to surface-fluency bias.
Phase 2 Analytical
Application of LREAD rubric;
eliciting rubric-scaffolded reasoning.
Measuring accuracy gains and
systematicity in detection.
Phase 3 Mastery
Detection on challenging domains using
internalized LREAD criteria.
Verifying domain-specific mastery
and expert-level consensus.
Table 3: Summary of the three-phase cognitive calibration framework, outlining the experimental design and evaluation targets for each calibration stage.
First, we incorporate a dedicated creativity dimension into the content category. Human writing frequently contains idiosyncratic metaphors, personal narratives, and non-statistical semantic leaps that depart from probabilistic expectations. In contrast, LLMs rely on the statistical recombination of training data, which often results in predictable and generic argumentation. We define creativity in this context as the presence of unique human signatures that violate the typical distribution of LLM outputs.
Second, we refine the scoring granularity to capture minute linguistic variances. Because modern LLMs achieve near-perfect grammatical accuracy, standard evaluation metrics often fail to differentiate synthetic text from human prose. Our granular system requires human experts to examine subtle patterns in lexical choice, morphological consistency, and sentence-level cohesion. By shifting the focus from broad fluency to specific structural regularities, we ensure that evaluators remain attentive to the subtle traces of machine intervention.
Table 2 summarizes the structured allocation of evaluation dimensions and corresponding point weights in the LREAD rubric. The rubric assigns a total of 100 points across content, organization, and expression, with expression accounting for the largest proportion. This asymmetric weighting reflects our empirical observation, derived from the Phase 1 blind study, that contemporary LLMs consistently achieve high scores on surface-level coherence and structural adequacy, while fine-grained expressive variation continues to serve as a robust indicator of human authorship.
Notably, the rubric evaluates creativity at both the content and expression levels. This dual placement enables the framework to capture human originality manifested through unconventional argumentative development as well as stylistic deviations from distributionally typical language patterns. By operationalizing creativity across multiple linguistic strata, the LREAD rubric equips annotators to identify subtle yet systematic traces of machine generation that remain undetected by automated detectors.
@@ 3.2 Three-Phase Cognitive Calibration
We propose a three-phase longitudinal calibration framework designed to trace and induce the evolution of human detection expertise. Rather than treating human judgment as static, this framework models detection as a learnable cognitive skill that emerges through iterative feedback and structured analytical scaffolding. The primary objective of the proposed framework is to systematically transform linguistically trained yet task-uncalibrated annotators into calibrated expert detectors, enabling them to reliably identify latent computational signatures that extend beyond surface-level linguistic fluency. The overall trajectory of this framework, detailing the transition of cognitive modes and evaluation targets at each stage, is summarized in Table 3.
Phase 1 establishes an intuitive baseline through a zero-shot blind detection task, as described in Section 2. During this phase, annotators rely solely on unstructured linguistic intuition, without access to a formal rubric or explicit decision criteria. We use this phase to quantify experts’ initial susceptibility to surface-fluency bias, whereby stylistic polish in LLM-generated text obscures underlying structural artifacts. Following Phase 1, we disclose the ground-truth labels and the identities of the four source models. This disclosure serves as an initial calibration signal, enabling annotators to retrospectively align their intuitive judgments with model-specific output patterns.
@@
Phase 3 represents domain-specific mastery. After a second disclosure of ground-truth labels, annotators perform a final blind detection task focused exclusively on elementary-level essays. We deliberately restrict this phase to the elementary domain, as it constitutes the most challenging setting for generation fidelity. We restrict this phase to the elementary domain, which poses the most stringent test of generation fidelity. Our evaluation set contains 10 essays: 2 human-written and 8 LLM-generated (2 each from GPT-4o, Solar, Qwen2, and Llama3.1), sampled using the same stratified strategy as in Phase 1. valentini-etal-2023-automatic show that state-of-the-art LLMs struggle to control lexical simplicity and readability when generating children’s stories, often exceeding age-appropriate vocabulary constraints. Complementing this finding, hassan2025evaluating demonstrate that expert evaluators systematically perceive LLM-generated child-like text as linguistically more advanced than the target age group. More broadly, wang-etal-2025-catch find that LLMs have difficulty imitating implicitly defined writing styles, tending to default to structurally over-regularized and stylistically polished outputs rather than capturing subtle, informal authorial signals. In the final phase, annotators apply the internalized rubric to detect persona-mismatch artifacts. We conclude by eliciting from calibrated experts explicit accounts of each source model’s linguistic fingerprints, alongside defining features of authentic human writing.

### v2

# From Intuition to Calibrated Judgment: A Rubric-Based Expert-Panel Study of Human Detection of LLM-Generated Korean Text

@@ Abstract
Distinguishing human-written Korean text from fluent LLM outputs remains difficult even for trained readers, who can over-trust surface well-formedness. We present LREAD, a Korean-specific instantiation of a rubric-based expert-calibration framework for human attribution of LLM-generated text. In a three-phase blind longitudinal study with three linguistically trained annotators, Phase 1 measures intuition-only attribution, Phase 2 introduces criterion-anchored scoring with explicit justifications, and Phase 3 evaluates a limited held-out elementary-persona subset. Majority-vote accuracy improves from 0.60 in Phase 1 to 0.90 in Phase 2, and reaches 10/10 on the limited Phase 3 subset (95% CI [0.692, 1.000]); agreement also increases from Fleiss’ $\kappa=-0.09$ to $0.82$. Error analysis suggests that calibration primarily reduces false negatives on AI essays rather than inducing generalized over-detection. We position LREAD as pilot evidence for within-panel calibration in a Korean argumentative-essay setting. These findings suggest that rubric-scaffolded human judgment can complement automated detectors by making attribution reasoning explicit, auditable, and adaptable.
From Intuition to Calibrated Judgment: A Rubric-Based Expert-Panel Study of Human Detection of LLM-Generated Korean Text
@@ 1 Introduction
The proliferation of large language models (LLMs) has reshaped digital communication and intensified concerns about academic (Perkins, 2023; Peng et al., 2023; Zhu et al., 2024) and creative integrity (Yang et al., 2024b; Wu et al., 2025). Recent models also generate text that is highly fluent and norm-conforming, which can appear authoritative even to trained readers. Although prior work reports systematic differences between human and machine writing, such as increased dependency depth in machine-generated outputs alongside greater semantic variety in human text (Zanotto and Aroyehun, 2025), these cues do not reliably translate into stable human judgments. In practice, human annotators often over-rely on surface well-formedness, as polished style and local grammaticality can obscure subtle structural irregularities, resulting in what we characterize as a persistent fluency trap for linguistically trained yet task-uncalibrated readers.
@@
A common expectation is that domain expertise should reduce this vulnerability. Evidence from English settings supports this view: frequent LLM users can outperform commercial systems such as GPTZero [^1] https://gptzero.me/ when identifying synthetic articles (Russell et al., 2025). At the same time, the majority of detection evidence and tooling remains English-centered, while high-stakes misuse increasingly occurs across diverse languages and writing norms. Non-English settings introduce additional sources of ambiguity, including language-specific morphology, orthography, and punctuation conventions, which can alter both the appearance of fluency and the distribution of artifacts (Park et al., 2025).
These observations motivate a different question: Is the primary bottleneck in human attribution a lack of linguistic knowledge, or the absence of an explicit diagnostic procedure? We investigate this question in Korean, a setting where spacing conventions, register variation, and punctuation optionality weaken the reliability of coarse fluency-based heuristics. In this sense, Korean constitutes more than a non-English extension; rather, it serves as a valuable test case for human-centered attribution.
To this end, we introduce LREAD, a Korean-specific rubric and calibration protocol for structured human attribution of LLM-generated text. Rather than treating human judgment as fixed, we study whether attribution accuracy, agreement, and auditability improve when linguistically trained annotators move from intuition-only screening to criterion-anchored analysis.
Our contributions are fourfold:
  • We present a three-phase expert-calibration protocol for studying human attribution as a learnable skill under strict information control.
  • We introduce LREAD as a Korean-specific instantiation grounded in national writing standards and adapted for attribution-oriented micro-diagnostics.
  • We provide pilot evidence, in a controlled argumentative-essay setting, that rubric scaffolding improves within-panel accuracy and agreement.
  • We release the full rubric, a calibrated cue taxonomy, and transparent reporting templates to support future work on auditable human-centered evaluation.
2 Related Work
Our work sits at the intersection of three lines of research: (i) automated LLM-generated text detection; (ii) linguistic profiling of LLM outputs; and (iii) human expertise in LLM-generated text detection.
2.1 Post-hoc Detection & LLM Watermarking
Prior work on attributing LLM-generated text has largely followed two directions: post-hoc detection (Mitchell et al., 2023; Su et al., 2023; Bao et al., 2024; Zhu et al., 2023; Yang et al., 2024a) and LLM watermarking (Kirchenbauer et al., 2023; Dathathri et al., 2024; Lu et al., 2024; Chen et al., 2024; Hou et al., 2024; Guo et al., 2024; Wang et al., 2025b; Park et al., 2026).
Post-hoc detection methods typically operate without altering the source model, making them suitable for retrospective auditing. DetectGPT (Mitchell et al., 2023) exploits probability curvature in the likelihood landscape, with subsequent work improving efficiency via log-rank–based estimators (Su et al., 2023) and conditional probability curvature (Bao et al., 2024). Other approaches leverage LLM querying through revision-based similarity analysis (Zhu et al., 2023) or analyze divergent continuations and n-gram distributions (Yang et al., 2024a). Despite strong empirical performance, these methods remain sensitive to domain shift and continued advances in generation quality.
LLM watermarking approaches address these limitations by embedding attribution signals directly into the decoding process. Kirchenbauer et al. (2023) introduce a foundational logit-bias watermarking framework, which subsequent work extends toward scalable deployment (Dathathri et al., 2024) and a range of design variants, including entropy-aware (Lu et al., 2024), lexical (Chen et al., 2024), semantic (Hou et al., 2024), and contextual (Guo et al., 2024) schemes. MorphMark (Wang et al., 2025b) advances this line by adaptively modulating watermark strength using the cumulative probability mass of green-list tokens to address the effectiveness–quality trade-off. WaterMod (Park et al., 2026) refines the paradigm through probability-ranked modular partitioning. Collectively, these approaches establish strong baselines while treating attribution as automated inference, with limited emphasis on human judgment.
2.2 Linguistic Profiling of AI-Generated Text
Zanotto and Aroyehun (2025) find that LLM-generated text often shows deeper and longer dependency structures than human writing, while human texts exhibit more diverse semantic content and greater overall variation. They further note that newer models become more human-like on syntactic depth and length metrics, shrinking earlier gaps. Within the Korean context, Park et al. (2025) identify language-specific artifacts in spacing, morphology, and punctuation that differentiate LLM-generated Korean text from human writing. These observations highlight the importance of attending to fine-grained linguistic realizations beyond surface-level fluency.
2.3 Human Expertise in AI Detection
Russell et al. (2025) demonstrate that frequent LLM users outperform many automatic detectors when identifying synthetic English articles, while providing interpretable rationales grounded in linguistic cues. Dugan et al. (2023) demonstrate that boundary detection remains challenging and exhibits substantial inter-annotator variability; nevertheless, targeted interventions are shown to produce measurable improvements over time.
@@ 3.1 Experimental Setup
We use the KatFish benchmark (Park et al., 2025) and focus on argumentative essays, which require a non-trivial balance of stance, evidence, and stylistic control. The evaluation set for each phase contains 30 essays selected through stratified sampling across elementary, middle, and high school levels. Each level contains two human-written essays and eight LLM-generated essays, with two samples each from GPT-4o, Solar-10.7B, Qwen2-72B, and Llama3.1-70B.
We recruit three advanced undergraduate annotators majoring in Korean Language and Literature. This results in a small yet linguistically trained panel characterized by a high degree of sensitivity to Korean grammar, discourse, and style, together with limited prior exposure to diverse LLM outputs. In Phase 1, annotators receive no rubric or explicit decision criteria and perform blind human-versus-AI attribution one essay at a time.
For a practical zero-shot judge baseline, we additionally evaluate GPT-5.2 Thinking, Gemini 3 Flash, and Claude Sonnet 4.5 under the same blind O/X instruction. We use separate generator and judge models to avoid self-recognition confounds and to estimate practical reviewing performance from strong independently developed LLMs.
@@
3.2 Human–LLM Detection Gap
Table 1 shows a substantial gap (36.67%p) between intuition-only human attribution and zero-shot LLM reviewing under identical instructions. On this benchmark, strong LLM judges achieve high accuracy, whereas linguistically trained humans are much less reliable when relying on intuition alone. We do not interpret this gap as evidence that human judgment is of limited use; rather, we take it to suggest that unguided human judgment is susceptible to a fluency trap, in which polished local well-formedness can obscure deeper attribution-relevant irregularities.
@@ 3.3 Analysis: Initial Human Perception
We summarize the heuristics that annotators use in the initial blind phase before they access gold labels or a formal rubric. We organize these heuristics around two recurring families: structural regularity and discourse plausibility. Annotators treat structural rigidity—overly template-like discourse organization with unnaturally “perfect” balance—and mechanical consistency—unusually uniform surface behavior in connective usage, punctuation, and spacing with little natural drift—as primary signals of machine generation.
They further characterize machine-generated essays as exhibiting overly literal adherence to prompts, limited semantic risk-taking, and forms of argumentation that remain normatively safe while lacking substantive depth. By contrast, they associate human writing with localized fallibility, creative deviation, and more organically grounded rhetorical progression. Importantly, these pre-calibration impressions are not yet stable diagnostic criteria; they are cue candidates that Phase 2 later re-anchors through explicit scoring rules.
@@ 4 LREAD: A Human-Centered Cognitive Calibration Framework
4.1 LREAD Construction from Linguistic Standards and Phase 1 Calibration
We derive LREAD from the 2025 national standards for argumentative writing proficiency issued by the National Institute of Korean Language [^2] Diagnostic Criteria for Argumentative Writing Ability (논증적 글쓰기 능력 진단 기준안), National Institute of Korean Language (국립국어원). ; however, the rubric itself is developed only after the completion of the Phase 1 blind evaluation. This sequencing is central to our study design. Once annotators finished the intuition-only task and received the gold labels, we mapped their reflections and disagreement cases onto criterion-anchored writing standards. Accordingly, LREAD is not a ready-made writing rubric; rather, it is a post-Phase-1 calibration instrument grounded in both external standards and empirically observed failure modes.
Importantly, the high-level rubric dimensions—Content, Organization, and Expression—remain stable, while some micro-diagnostics are intentionally language- and genre-sensitive. In the present Korean instantiation, these include orthography, spacing, register control, and natural Korean phrasing judgments that are difficult to capture through coarse notions of fluency alone. This design preserves comparability with established writing assessment principles while increasing sensitivity to the specific artifacts that emerged in our blind human–LLM discrimination task.
We make two detection-oriented adaptations. First, we retain creativity-related criteria because human essays often depart from probabilistically typical realization through localized novelty, idiosyncratic phrasing, or non-formulaic argumentative development. Second, we increase scoring granularity so that annotators attend to subtle lexical, orthographic, register, and cohesion-level variation that modern LLMs often smooth out. The resulting rubric is designed to surface cases in which machine-generated essays appear systematically too correct, too regular, or too register-stable relative to authentic student writing.
Table 2 summarizes the structured allocation of evaluation dimensions and point weights in LREAD. The rubric assigns 100 total points across content, organization, and expression, with the largest share placed on expression-level criteria. This weighting reflects our empirical observation from the Phase 1 blind study: contemporary LLM outputs often satisfy broad coherence and structural adequacy, while finer-grained expressive variation remains more diagnostic for human authorship. Creativity is therefore included as one informative component of the rubric rather than as its sole or primary discriminator; specifically, it contributes 20 of the 100 total points across content and expression.
By combining criterion-anchored writing standards with post-hoc calibration from real annotator errors, LREAD operationalizes human-sensitive cues across multiple linguistic strata. This enables annotators to move beyond surface-level fluency and identify subtle yet systematic traces of machine generation that frequently elude automated detection.
@@ 4.1 LREAD Construction from Linguistic Standards and Phase 1 Calibration
Phase 1 Intuitive baseline Blind attribution using unstructured linguistic intuition Quantifying susceptibility to the fluency trap and eliciting early cue candidates
Phase 2 Criterion-anchored analysis Application of the LREAD rubric with explicit scoring and justification Measuring gains in accuracy, consistency, and reasoning transparency
Phase 3 Internalized application Blind attribution on a limited held-out elementary-persona subset using internalized LREAD criteria Testing application to a residual difficulty condition and examining consensus on a small held-out set
Table 3: Summary of the three-phase calibration framework. The protocol is designed to study human attribution as a learnable, auditable procedure rather than a fixed trait.
@@ 4.2 Three-Phase Cognitive Calibration
We propose a three-phase longitudinal calibration framework designed to study whether human attribution can improve through explicit feedback and structured analytic scaffolding. Rather than treating human judgment as static, the framework models attribution as a learnable procedure whose quality can be assessed through changes in accuracy, agreement, and reasoning transparency. The overall trajectory of this framework, including the transition of cognitive modes and evaluation targets at each stage, is summarized in Table 3.
Phase 1 establishes an intuitive baseline through a zero-shot blind detection task, as described in Section 3. During this phase, annotators rely solely on unstructured linguistic intuition, without access to a formal rubric or explicit decision criteria. We use this phase to probe initial failure modes, including susceptibility to surface-fluency bias, whereby stylistic polish in LLM-generated text can obscure underlying structural artifacts. Following Phase 1, we disclose the ground-truth labels and the identities of the four source models. This disclosure serves as an initial calibration signal, enabling annotators to retrospectively align their intuitive judgments with recurring output patterns and disagreement cases.
@@
Phase 3 evaluates internalized application on a limited held-out subset. After a second disclosure of ground-truth labels, annotators perform a final blind detection task focused exclusively on elementary-level essays. Our focus on the elementary-persona condition does not stem from the assumption that it constitutes the most difficult setting in all cases; rather, it reflects the fact that this condition remained the primary residual challenge after Phase 2 calibration and therefore serves as a focused evaluation of whether the rubric can be applied without additional external scaffolding. Our evaluation set contains 10 previously unseen essays: 2 human-written and 8 LLM-generated (2 each from GPT-4o, Solar, Qwen2, and Llama3.1), sampled using the same stratified strategy as in Phase 1.
We treat this phase as a test of whether annotators can detect persona-mismatch artifacts, that is, discrepancies between the assumed student persona and the realized linguistic or discourse properties of the text. Such artifacts include essays framed as the work of an elementary student but marked by overly mature vocabulary, adult-like institutional reasoning, or an unusually polished and textbook-like organization of discourse. This framing is consistent with prior work showing that state-of-the-art LLMs struggle to control age-appropriate lexical simplicity and readability in child-oriented generation (Valentini et al., 2023), and that expert evaluators often perceive LLM-generated child-like text as linguistically more advanced than the target age group (Hassan et al., 2025). More broadly, Wang et al. (2025a) find that LLMs have difficulty imitating implicitly defined writing styles, often defaulting to structurally over-regularized and stylistically polished outputs rather than capturing subtle, informal authorial signals.
In the final phase, annotators apply the internalized rubric in a blind binary human-versus-AI attribution task, with optional uncertainty marking. Only after completing the blind task do we elicit explicit accounts of each source model’s linguistic fingerprints, alongside defining features of authentic human writing; these post-hoc reflections are used solely for qualitative analysis. Because the held-out subset is limited in size, we interpret Phase 3 as a preliminary assessment of criterion internalization rather than a definitive demonstration of robust generalization across domains.